OpenStudy (anonymous):

simplify: 2x - 1 ---- ---- x^2-9 x-3 a) 2x-1 ------ x^2-9 b) 2x-1 ------ x+3 c) 2x-1 ------ x-3 d) 1 ---- x+3 e) 1 ---- x-3

OpenStudy (anonymous):

The answer is d)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

2x/(x-3)(x+3) - 1/x-3 x-3/(x-3)(x+3) 1/x+3

jimthompson5910 (jim_thompson5910):

\[\frac{2x}{x^2-9}-\frac{1}{x-3}\] \[\frac{2x}{(x-3)(x+3)}-\frac{1}{x-3}\] \[\frac{2x}{(x-3)(x+3)}-\left(\frac{1}{x-3}\right)\left(\frac{x+3}{x+3}\right)\] \[\frac{2x}{(x-3)(x+3)}-\frac{1(x+3)}{(x-3)(x+3)}\] \[\frac{2x}{(x-3)(x+3)}-\frac{x+3}{(x-3)(x+3)}\] \[\frac{2x-(x+3)}{(x-3)(x+3)}\] \[\frac{2x-x-3}{(x-3)(x+3)}\] \[\frac{x-3}{(x-3)(x+3)}\] \[\frac{1}{x+3}\] So \[\frac{2x}{x^2-9}-\frac{1}{x-3}=\frac{1}{x+3}\]